# AdagradLSPIA: Integrating adaptive optimization into least squares progressive iterative approximation

### v1

@@ 3 Experimental study
This section details the implementation of the AdagradLSPIA method and presents the experimental results of its comparison with the standard LSPIA method in the context of B-spline surface fitting.
@@ 3.1 Implementation details
The experiments utilize a topologically rectangular set of $251\times 251$ data points representing an automobile hood that features a combination of smooth regions, sharp transitions, and subtle design details (see Figure 1). Parameter values for the data points were assigned using a uniform parameterization method in both parametric directions.
A bicubic tensor product B-spline surface with $50\times 50$ control points was employed for fitting. The knot vectors to define the B-spline basis functions were constructed using the averaging technique (see [14, 15]). For initialization, all control points were set to zero.
Fig. 1: Data points representing an automobile hood model.
The fitting error was computed at each iteration as the sum of the squared Euclidean distances between the data points and their corresponding points on the fitting surface. The iterative process terminated either when the absolute difference between consecutive fitting errors was less than $10^{-7}$ or when the maximum of $10^{3}$ iterations was reached.
@@ 3.2 Results and discussion
The initial set of experiments investigated how global weight affected the performance of both methods. LSPIA was tested with 200 global weight values sampled uniformly from the interval $(0,0.2]$, while AdagradLSPIA was examined over $(0,20]$. The results (Figure 2) show that LSPIA converged only within a limited range of global weights; outside this range, it reached the maximum iteration limit without achieving the desired accuracy. In contrast, AdagradLSPIA converged consistently across the entire tested range, demonstrating its robustness to variations in global weight.
@@
From the experiments, global weight values yielding minimal fitting error, minimal elapsed time, and minimal iterations were identified for both methods. Table 1 summarizes these results, showing that AdagradLSPIA achieved better accuracy, reduced computational time, and required fewer iterations compared to LSPIA. For example, the minimal fitting error for AdagradLSPIA was $8.75530\times 10^{-7}$, which is significantly lower than $1.78478\times 10^{-6}$ achieved by LSPIA.
@@
The final fitting surfaces (Figure 4), obtained using the global weights corresponding to the minimal fitting errors, and their reflection lines (Figure 5) provide a qualitative comparison of the results. The reflection lines show the ability of AdagradLSPIA to capture geometric details and maintain surface smoothness.
Fig. 2: Comparison of the LSPIA and AdagradLSPIA methods across a range of global weight values. The plots depict the variation of fitting error, elapsed time (in s), and the number of iterations required to converge, highlighting the broader convergence region and superior performance of AdagradLSPIA.
@@
LSPIA 0.016754 3.47095E-06 0.398633 153
AdagradLSPIA 16.200000 1.15355E-06 0.145413 56
Fig. 3: Convergence plots for the LSPIA and AdagradLSPIA methods using optimized global weight values. The plots show the reduction in fitting error over elapsed time and iterations, demonstrating the faster convergence of AdagradLSPIA. (For interpretation of the references to color in this figure legend, the reader is referred to the web version of this article.)
Fig. 4: The final fitting surfaces generated by the LSPIA and AdagradLSPIA methods. The surfaces were obtained using the respective global weights that minimize the fitting error.
Fig. 5: Reflection lines (zebra mappings) on the fitting surfaces generated by the LSPIA and AdagradLSPIA methods. These lines demonstrate the ability of AdagradLSPIA to capture geometric details and maintain surface smoothness.
@@ 3.3 Summary of results
The experiments conclusively demonstrate the advantages of AdagradLSPIA over the standard LSPIA method. AdagradLSPIA consistently achieved higher accuracy, faster convergence, and improved surface quality. Its robustness to global weight variations and adaptability to the geometric characteristics of the data make it a superior choice for surface fitting tasks.
@@ 4 Concluding remarks
This paper introduced the Adaptive Gradient Least Squares Progressive Iterative Approximation (AdagradLSPIA) method, a significant enhancement of the standard LSPIA technique for surface fitting. By integrating the Adagrad optimization algorithm into the geometric iterative framework, AdagradLSPIA dynamically adjusts weights based on accumulated gradient information, resulting in more stable and efficient convergence.
Theoretical analysis showed that AdagradLSPIA maintains the guaranteed convergence of LSPIA while achieving a sublinear convergence rate of $\mathcal{O}(1/\sqrt{K})$. The experimental results confirmed the practical advantages of the method, including faster convergence, reduced computational cost, and improved accuracy compared to the standard approach. In particular, AdagradLSPIA exhibited superior performance in a range of global weight values and consistently produced smooth and accurate fitting surfaces, as validated by reflection line analysis.
In addition to its promising results, AdagradLSPIA opens avenues for further research. The integration of other advanced adaptive optimizers, such as Adam [16], Adadelta [17], Nadam [18], hypergradient descent [19], or AMSGrad [20], could potentially improve the efficiency and robustness of the method. Future work may also extend the method to handle more complex geometries. Overall, AdagradLSPIA represents a robust and efficient tool for surface fitting, with significant potential for applications in geometric modeling, CAD, and reverse engineering.

### v2

@@ 3 Experimental study
This section details the implementation of the AdagradLSPIA method and presents the experimental results of its comparison with the standard LSPIA method in the context of tensor product B-spline surface fitting. The experiments were conducted on both noise-free and noisy data to examine the performance and robustness of the proposed method.
@@ 3.1 Implementation details
The experiments utilize a topologically rectangular set of $251\times 251$ data points representing an automobile hood that features a combination of smooth regions, sharp transitions, and subtle design details (see Figure 1). For the noisy data case, zero-mean Gaussian noise with a standard deviation of $0.02$ was added to the original data. Parameter values for the data points were assigned using a uniform parameterization method in both parametric directions.
Fig. 1: Data points representing an automobile hood model, including noise-free data (left) and noisy data (right) with added zero-mean Gaussian noise (standard deviation 0.02).
A bicubic tensor product B-spline surface with $25\times 25$ control points was used for the fitting. The knot vectors to define the B-spline basis functions were constructed using the averaging technique (see [19, 20]). For initialization, all control points were set to zero.
The fitting error was computed at each iteration as the sum of the squared Euclidean distances between the data points and their corresponding points on the fitting surface. The iterative process terminated either when the absolute difference between consecutive fitting errors was less than the prescribed tolerance ($10^{-7}$ for noise-free data and $10^{-7}$ for noisy data) or when the maximum of $10^{3}$ iterations was reached.
@@ 3.2 Results and discussion
The experiments analyzed the performance of LSPIA and AdagradLSPIA on noise-free and noisy data by examining their sensitivity to the global weight, convergence behavior, and final surface quality.
@@
The initial set of experiments investigated how the global weight affected the performance of the LSPIA and AdagradLSPIA methods. LSPIA was tested with 200 global weight values sampled uniformly from the interval $(0,0.02]$, while AdagradLSPIA was examined over $(0,20]$. The results (Figure 2 and Figure 3) show that LSPIA converged only within a limited range of global weights; outside this range, it reached the maximum iteration limit without achieving the desired accuracy. In contrast, AdagradLSPIA converged consistently throughout the tested range, demonstrating its robustness to variations in global weight.
@@
From the experiments, global weight values yielding minimal fitting error, minimal elapsed time, and minimal iterations were identified for both methods. Table 1 summarizes these results, showing that AdagradLSPIA achieved better accuracy, reduced computational time, and required fewer iterations compared to the LSPIA method. For example, the minimal fitting error for AdagradLSPIA applied to noise-free data was $8.75530\times 10^{-7}$, which is significantly lower than $1.78478\times 10^{-6}$ achieved by LSPIA.
@@
The final fitting surfaces (Figure 5), obtained using the global weights corresponding to the minimal fitting errors, and their reflection lines (Figure 6) provide a qualitative comparison of the results. The reflection lines show the ability of AdagradLSPIA to capture geometric details and maintain surface smoothnes, even on noisy data.
Fig. 2: Comparison of the LSPIA and AdagradLSPIA methods for noise-free data across a range of global weight values. The plots depict the variations in fitting error, elapsed time (in s), and the number of iterations required to converge, highlighting the broader convergence region and superior performance of AdagradLSPIA.
Fig. 3: Comparison of the LSPIA and AdagradLSPIA methods for noisy data across a range of global weight values. The plots depict the variations in fitting error, elapsed time (in s), and the number of iterations required to converge, highlighting the broader convergence region and superior performance of AdagradLSPIA.
@@
LSPIA 0.016754 3.47095E-06 0.325270 153
AdagradLSPIA 16.200000 1.15355E-06 0.122095 56
Noiy data
Minimal fitting error
LSPIA 0.016467 3.27684E-02 0.144296 61
AdagradLSPIA 2.900000 3.04729E-02 0.025656 12
Minimal elapsed time
LSPIA 0.015226 3.74664E-02 0.070095 36
AdagradLSPIA 1.100000 3.20504E-02 0.014523 8
Minimal number of iterations
LSPIA 0.015035 3.78637E-02 0.070474 36
AdagradLSPIA 1.100000 3.20504E-02 0.014523 8
Fig. 4: Convergence plots for the LSPIA and AdagradLSPIA methods using optimized global weight values. The plots show the reduction in fitting error over elapsed time and iterations for both noise-free and noisy data, demonstrating the faster convergence of AdagradLSPIA. (For interpretation of the references to color in this figure legend, the reader is referred to the web version of this article.)
Fig. 5: The final fitting surfaces generated by the LSPIA and AdagradLSPIA methods for noise-free and noisy data. The surfaces were obtained using global weights that minimize the fitting error (Table 1).
Fig. 6: Reflection lines (zebra mappings) on the fitting surfaces generated by the LSPIA and AdagradLSPIA methods for noise-free and noisy data. These lines demonstrate the ability of AdagradLSPIA to capture geometric details and maintain surface smoothness, even on noisy data.
@@ 3.3 Summary of results
The experiments demonstrate the advantages of AdagradLSPIA over the standard LSPIA method. AdagradLSPIA consistently achieved higher accuracy, faster convergence, and improved surface quality. Its robustness to global weight variations and adaptability to the geometric characteristics of the data make it a superior choice for surface fitting tasks.
@@ 4 Concluding remarks
This paper introduced the Adaptive Gradient Least Squares Progressive Iterative Approximation (AdagradLSPIA) method, a significant enhancement of the standard LSPIA technique. By integrating the Adagrad optimization algorithm into the geometric iterative framework, AdagradLSPIA dynamically adjusts weights based on historical gradient information, resulting in more stable and efficient convergence.
The experimental results confirmed the practical advantages of the AdagradLSPIA method, including faster convergence, reduced computational cost, and improved accuracy compared to the standard approach. In particular, AdagradLSPIA exhibited superior performance in a range of global weight values and consistently produced smooth and accurate fitting surfaces, as validated by reflection line analysis.
The computational complexity of AdagradLSPIA remains consistent with the standard LSPIA method, with the adaptive weighting mechanism introducing minimal additional computational overhead. Compared to other enhanced LSPIA variants, it offers an efficient and practical solution for surface fitting tasks, especially when faster convergence and robustness are required.
In addition to its promising results, AdagradLSPIA opens avenues for further research. The integration of other advanced adaptive optimizers, such as Adam [21], Adadelta [22], Nadam [23], hypergradient descent [24], or AMSGrad [25], could potentially improve the efficiency and robustness of the method. Future work may also extend the method to handle more complex geometries.
Overall, AdagradLSPIA represents a robust and efficient tool for surface fitting, with significant potential for applications in geometric modeling, CAD, and reverse engineering.